What is the atomic mass of hafnium if out of every 100 atoms, 5 have mass 176, 19 have mass 177, 27 have mass 178, 14 have mass 179, and 35 have mass 180.0?


1
Expert's answer
2020-11-04T14:24:14-0500

Atomic mass of hafnium= "\frac{5}{100}×176 + \frac{19}{100}×177 + \frac{27}{100}×178 + \frac{14}{100}×179 + \frac{35}{100}×180"

Atomic mass of hafnium= 178.55